OpenStudy (anonymous):

explain what 4^3/2 means

OpenStudy (anonymous):

Do you mean: \[\frac{4^3}{2}\] or \[4^(2/3)\]

OpenStudy (anonymous):

\[4^{3/2}\]

OpenStudy (anonymous):

The square root of 4 cubed, is this sufficient explanation?

OpenStudy (anonymous):

Yes can you calculate Calculate 4^3/2 in two different ways.

OpenStudy (anonymous):

I think I answered it, either cube 4 and take the square root or take the square root of 4 and cube it.

OpenStudy (anonymous):

Is it clear?

OpenStudy (anonymous):

ya i understand what you are saying i am not sure what my instructor is asking

OpenStudy (anonymous):

He asked you to calculate it in two different ways?

OpenStudy (anonymous):

sqrt(4*4*4) = sqrt(64) = 8 sqrt(4) * sqrt(4) * sqrt(4) = 2*2*2 = 8